How do you solve this equation? (-6+4x=x+9)
Answer by skue(3)

In solving such equations, you combine similar terms.
-6 + 4x = x + 9 ( so, transpose x to the other side)

-6 + 4x - x = 9. » NOTE: when transposng variables, their signs always changes
Then, transpose -6 to the other side, the equation then bcomes
4x - x = 9 + 6.
Perform the indicated operations
4x - x = 9 + 6
3x = 15 (divide both sides by 3 to give you x or in other words, rem0ve the coefficient of x)
x = 5